Hello,
This math problem I need some assitance with. I had been thinking that I have to convert the 2 equations into y=mx+b form..can you please help me out? Because I am not sure how to graph them. Thanks.
Graph the following equations. Make sure you clearly show the x and y intercepts on your graph.
Graph:
(2/3) x = 4 + y

Graph:
-3 y + 12x = 15

Answer by ptaylor(2198)

You don't need to convert them into standard form (y=mx+b)
The easiest way to get two points on the straight line is to first let x=0, then solve for y and next let y=0 then solve for x. That will give you two points as well as your x and y intercepts. I'll do the first one and you can do the second one.
(2/3)x=4+y
Let x=0
0=4+y
y=-4
P1(0,-4)
Let y=0
(2/3)x=4+0
(2/3)x=4 multiply each side by (3/2)
x=6
P2(6,0)
Two points determine a straight line, so there you have 'em.
Hope this helps---ptaylor